OpenStudy (anonymous):

Type the correct answer in the box. Use numerals instead of words. If necessary, use / for the fraction bar(s). The circumference of a pizza is approximately 56.52 inches. One-fourth of the pizza has been eaten. Enter the area of the remaining pizza, in square inches. Round your answer to the nearest hundredth. Use 3.14 for π.

Directrix (directrix):

>>****circumference****** of a pizza is approximately 56.52 inches @FaiqRaees

OpenStudy (anonymous):

I am confused on how to figure it out

OpenStudy (faiqraees):

Circumference = 2πr 56.52 = 2πr r = 56.52/2π 3/4 of pizza area = 3/4 x πr² Substitute the r here and you will have your answer

Directrix (directrix):

@breihl What is the value of this: 2 * (3.14) = ?

Directrix (directrix):

That is not the answer to the question but a step towards the answer.